Give the component form of the resultant vector in the following.
NOTE: Answer must be typed in using the following format -- including the parentheses: (#,#)
u = (16, -8)
v = (-5, 3)
u + v = ?

Answer by jim_thompson5910(35256)

Add up the corresponding components

u + v = (16, -8) + (-5, 3)

u + v = (16 + (-5), -8+3)

u + v = (11, -5)